OpenStudy (anonymous):

what property does 0+153 show. I will give medal

OpenStudy (123qwet12):

That's not a complete equation?

OpenStudy (anonymous):

the whole equation is Indicate which property is illustrated in Step 4. Step 1 (-3) + 3 + 9(11 + 6) + = (-3) + 3 + (9 • 11) + (9 • 6) Step 2 = (-3) + 3 + (99) + (54) Step 3 = (-3) + 3 + 153 Step 4 = 0 + 153 Step 5 = 153 A. inverse property of addition B. inverse property of multiplication C. associative property of addition D. commutative property of division

OpenStudy (123qwet12):

Well it's a complete equation, but I've never seen a property like that

OpenStudy (123qwet12):

Oh well you have to put that up too because we can't see that.

OpenStudy (anonymous):

sorry its just i put it up and no one replied

OpenStudy (123qwet12):

inverse property of addition